If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
thank you for all your help
: If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
thank you for all your help

Answer by ptaylor(1326)



Let x=amount of time it takes them working together to do the job
Sam works at the rate of 1/4 job per day
Lisa works at the rate of 1/6 job per day
Tom works at the rate of 1/2 job per day
Together tey work at the rate of 1/4 +1/6 +1/2 job per day
1/4 +1/6 +1/2=3/12 +2/12 +6/12=11/12 job per day
So, our equation t solve is:
(11/12)*x=1 (1 job, that is) multiply each side by 12
11x=12 divide each side by 11
x=1 1/11 days----------time it takes them all working together

Hope this helps---ptaylor